Okay, I have to find the area of a regular polygon.
This question is:
Find the area of a pentago with a side of 10 ft.
I keep getting 216.5, but the answer is 172. How did I mess up?

Found 2 solutions by solver91311, rfer:
Answer by solver91311(24713)



I don't know how you "messed up" because you didn't share any of your calculations, but mess up you certainly did.

To find the area of a regular pentagon, square the side length and multiply by the factor




John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by rfer(16322)

Use the equation.
A=1.72048*s^2
A=1.72048*10^2
A=1.72048*100
A=172.048
round off to
A=172
This works for a regular pentagon